Given f(x)=5x+2 and g(x)=2x+1 find (f-g)(-1) and (fg)(2)
Answer by edjones(7311)

f(x)=5x+2 and g(x)=2x+1
(f-g)(-1)=((5*-1)+2)-((2*-1)+1)
=(-5+2)-(-2+1)
=-3+2-1
=-2
.
fg(2)=(5*2+2)(2*2+1)
=12*5
=60
.
Ed